OpenStudy (anonymous):

1. Let a and b be equal non-zero quantities a = b 2. Multiply through by a a^2 = ab 3. Subtract b^2 a^2 - b^2 = ab - b^2 4. Factor both sides (a - b)(a + b) = b(a - b) 5. Divide out (a - b) a + b = b 6. Observing that a = b b + b = b 7. Combine like terms on the left 2b = b 8. Divide by the non-zero b 2 = 1

OpenStudy (anonymous):

"a = b"\[ a-b=0 \] "5. Divide out (a - b)" Divide out \(0\) No

OpenStudy (anonymous):

got it!ty

OpenStudy (anonymous):

\[ (a - b)(a + b) = b(a - b) \implies \frac{a - b}{a-b}(a + b) = b \] You say \(0/0=1\), but I say: \(0/0=1/2\)

OpenStudy (anonymous):

\[ \frac 12(a+b)=b\\ \frac 12(b+b)=b\\ b=b \]And I'm right.

OpenStudy (anonymous):

\[ 0/0 \]Is an indeterminate form. In this case we can determine that is is \(1/2\).

OpenStudy (anonymous):

ty